OpenStudy (anonymous):

anti-derivative of (4th root of x^3 +1)dx

OpenStudy (amistre64):

cant do it without an x^2 component i beleive

myininaya (myininaya):

\[\int\limits_{}^{}(\sqrt[4]{x^3}+1) dx\]

myininaya (myininaya):

is that right?

OpenStudy (amistre64):

well, myins is doable lol

OpenStudy (anonymous):

yes that is correct.

myininaya (myininaya):

\[\int\limits_{}^{}(x^\frac{3}{4}+1)dx\]

myininaya (myininaya):

\[\frac{x^{\frac{3}{4}+1}}{\frac{3}{4}+1}+x+C\]

myininaya (myininaya):

\[\int\limits_{}^{}x^n dx=\frac{x^{n+1}}{n+1}, n \neq -1\]